Find the capacitance of a parallel plate capacitor having plates of area 1.00 m2 that are separated by 0.400 mm of Teflon.
1     
Answer #1

k=2.1

so, C=keo*A/d

=2.1*8.85*10^-12*1/(4*10^-4)

=4.646*10^-8 F

=0.004646 uF